# Discrete Chi-Square Method can model and forecast complex time series, like El Nino data between 1870 and 2024

### 04 Mar 2026 v7

@@ Abstract
Forecasting El Niño is one of the greatest challenges of science. We show how intensive, large and accurate time series allow us to see through time. Our Discrete Chi-square Method (DCM) can detect arbitrary trend and signal(-s) combinations. It can forecast complex time series. The widely-used Discrete Fourier Transform (DFT) and other frequency-domain parametric time series analysis methods have many application limitations. None of those limitations constrains the DCM. Our simulated time series analyses ascertain the revolutionary Window Dimension Effect (WD-effect): “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.” The simulations also expose the DFT’s weaknesses and the DCM’s efficiency. The DCM’s backbone is the Gauß-Markov theorem that the Least Squares (LS) is the best unbiased estimator for linear regression models. DCM can not fail because this simple method is based on the computation of a massive number of linear model LS fits. The Fisher-test gives the signal significance estimates and identifies the best DCM model from all alternative tested DCM models. The analytical solution for the non-linear DCM model is an ill-posed problem. We present a computational well-posed solution. The best DCM model must be correct if it passes our Forecast-test. Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series. In our appendix, we show that the DCM can model and forecast El Nino data between 1870 and 2024. An immediate, independent and objective validity check of our analysis may save some money.
@@ 1 Introduction
This study proceeds through following stages. We present the DCM formulation (Section 2.1). The WD-effect is discussed (Section 2.2). Complex time series are simulated using the DCM model (Sect. 2.3). The DFT and its ALs are presented (Section 2.4). We compare our DCM to the renowned DFT. The time series analysis of simulated data sets shows that the ALs of DFT do not constrain DCM (Sections 3.1-3.8). Then, we present the identification of the best DCM model for the data (Section 3.9), the DCM significance estimates (Section 3.10), the well-posed computational DCM model solution (Section 3.11) and the DCM forecasts (Section 3.12). Finally, we discuss our results (Section 4). In our Appendix, we show that the DCM can model and forecast El Niño (Section A1).
@@ 2.1 Discrete Chi-square method (DCM)
The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [31, TSPA, Equation 1] and the Continuous Period Search [39, CPS, Equation 3]. TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$. TSPA can detect pure sine and double wave signals ($K_{2}=1$ or 2). The extension of TSPA, the CPS method, tests three alternatives: $K_{2}=1$ or 2, or no signal at all. The DCM model is more sophisticated because it can have any arbitrary $K_{1}$, $K_{2}$ and $K_{3}$ combination. The DCM sum $g(t)$ =h(t)+p(t)=h(t)+p(t) of arbitrary periodic and aperiodic functions represents a universal model because the innumerable $K_{1}$, $K_{2}$ and $K_{3}$ combinations allow unlimited complexity. Even more complex signal shape $(K_{2})$ combinations could be used. For example, these shapes could be $K_{2,i=1}=3$ (3rd harmonic), $K_{2,i=2}=1$ (pure sine) and $K_{2,i=3}=2$ (double wave), where if $K_{2,i}$ denotes the shape of $i$:th signal.
@@ 2.4 Discrete Fourier Transform (DFT)
The DFT is one of the most prevalent frequency-domain parametric time series analysis methods for unevenly spaced time series. It searches for the best pure sine model for the data. The equivalent DCM model has the orders $K_{1}=1,K_{2}=1$ and $K_{3}=0$. Any time series analysis method must be remarkable if it performs better than the distinguished DFT. Therefore, we will compare how DCM and DFT perform in the analyses of simulated time series (Sections 3.1-3.7). We search for signals in these simulated time series by using the frequently applied DFT version formulated by [24] [^1] This paper had over 2600 citations in December 2025. Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [24, Equation 1]. The notations for the DFT model are
@@
The time series is “too short” [32, AL6]
@@
The time series is non-stationary because it contains “a trend” [45, AL7]
@@
Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [32, AL9].
@@ 3.9 Best model
In the above example, the Fisher-test finds the correct number of signals $(K_{1})$ and the correct trend $(K_{3})$ for the pure sine signal alternative $(K_{2}=1)$. We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large. One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu [30, Table S7].
@@ 3.10 Significance estimates
After the detection of the first strongest signal, the Fisher-test critical level $Q_{\mathrm{F}}$ values increase for the next detected weaker signals. In other words, the signal significances decrease. Typical examples can be found in Jetsu [30, Tables S5-S1]. No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected. For the Model 1 simulated $n=100$ and ${\mathrm{SN}}=100$ combination, Table 8 shows that this time series contains only one signal.
For this Model 1 time series, there is no need to discuss the DFT significance estimates [24, their Equation 22] because this method fails to detect the correct period.
@@ 3.11.3 Stability ( $\mathrm{C}_{3}$)
The artificial bootstrap data sets (Equation 31) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation). We routinely check the stability of these solutions [30, Figure S5]. The unstable models, where the model parameter changes are large, are rejected (Section 2.1: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}” models).
@@ 3.12 Forecast
There are numerous techniques for forecasting a time series [e.g., 23, 28, 33]. The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast. We divide all data into the forecasting data and the forecasted data. The time points, the observations and the errors of these samples are
@@ Results
We use the first half of all yearly data $(n=78)$ to forecast the second half $(n=77)$. The results of weighted DCM analysis are given in Table S2. The periods of the three strongest signals are $P_{1}=5.580\pm 0.085$, $P_{2}=12.82\pm 0.40$ and $P_{3}=19.30\pm 0.83$ years (Model $\mathcal{M}$=3). The forecasts of the two and three signals models, open circles, are shown in Figures S1 and S2. For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signal models are 421, 336, 370 and 428, respectively. Hence, the two signal $\mathcal{M}$=2 model gives the best “Big Wave” forecast.
@@
For all weighted data, the detection of the first “Big Wave” signal is absolutely certain. The Fisher-test comparison between the $K_{3}=0$ pure noise model $g_{0,1,0,\chi^{2}}$ and the one signal model $g_{1,1,1,\chi^{2}}$ gives signal significance $Q_{\mathrm{F}}=3.9\times 10^{-13}$ (Table S6, $\mathcal{M}$=1). All other polynomial model alternatives, $1\leq K_{3}\leq 7$, are also rejected. The second and third “Big Wave” signal significances are $Q_{\mathrm{F}}=2.7\times 10^{-7}$ and $Q_{\mathrm{F}}=0.00038$ (Table S3: $\mathcal{M}$=2-3).
@@
Our mathematical model can detect regularities that the current physical models fail to detect, let alone forecast. Mathematics detects signals that should not be there according to Physics. The counterargument is that simple mathematics can not explain a complex physical phenomenon. We argue that the El Niño physical model for the “Big Wave” is simple. The solar energy output changes could determine the “Big Wave”. It is no surprise that the Sun warms the largest battery on the Earth, the Pacific Ocean. And if many cycles moderate the solar energy output, it is no surprise either that we have failed to discover the interference of these cycles, the “Big Wave”, in our yearly climate temperatures. The immense Pacific Ocean surface acts like a giant thermometer that measures solar energy output in real time.
@@
Note added to the proof
We arrive at a "Mexican standoff" because only future data can confirm that our “Big Wave” forecasts are correct. Fortunately, one minor detail saves us one year of waiting.
On January 29th, 2026, we computed the yearly values from the NOAA1 data. Four monthly values for the year 2025 were missing and our analysed data ended in the year 2024. Today February 25th 2026, NOAA1 still has these same four missing monthly values for the year 2025 (“-99.99”). To get these missing values, we are advised to use the Climate Prediction Center (CPC) data in NOAA2.666https://www.cpc-ncep.noaa.gov/products/
@@
All twelve monthly values for the year 2025 are given in Table S7 (column 2). These values give $y_{i}\pm 1\sigma_{i}=-0.26\pm 0.21$ $\mathrm{C}^{o}$. A red circle now highlights this observed new 2025 year value which is added into Figures S3e and Figures S4e. Our “Big Wave” forecast for the year 2025 succeeds. The year 2025 forecast for the best $\mathcal{M}$=3 model for all data (Table S3) is more accurate than the forecast for the unstable $\mathcal{M}$=4 model. Our forecast for this 2025 observation can, of course, be dismissed just as one lucky guess. Who wants to take the risk that our severe “Big Wave” forecasts for the years 2030 and 2031 are also “lucky”?
We also analyse the monthly SST values (Tables S8-S19, Figures S5-S16). The red dots in our figures highlight the available new monthly values between January 2025 and February 2026. Most of those red dots hit their mark. We consistently detect the “new” 3.6 and 4.6 year periods from these monthly values. The monthly mean computation eliminates the detection of these periods from yearly data. We conclude that the chaotic $\pm 1$$\mathrm{C}^{o}$ weather changes can prevent accurate DCM monthly forecasts. However, we can compute the “Big Wave” monthly values (Table S20). We admit that this does not fully solve the El Niño riddle but it may help. If no internal mechanism can cause the strictly multi-periodic “Big Wave” in the Pacific Ocean then the Sun is the only possible external cause. The planets may indeed cause regular solar cycles [30].
@@ Acknowledgements
All data files and Python codes are stored to
@@
Figure S3: “Big Wave” for all yearly mean data model $\mathcal{M}$=3 in Table S3. Red circle is added observed 2025 year value. Otherwise as in Figure S1.
@@
Table S2: Periods in first half yaerly mean data. Otherwise as in Table S1.
@@
Table S4: “Big Wave”yearly forecasts (Figures S3d and S4d: open circles).
@@
Table S7: Year 2025.
Temperature anomalies $(y)$ of 12 months.
(1) NOAA1
(2) NOAA2
$y$ $y$
$\mathrm{C}^{o}$ $\mathrm{C}^{o}$
@@
-0.50 -0.4
-0.36 -0.2
-0.13 -0.1
-0.09 0.0
0.02 0.0
0.14 0.0
0.05 -0.1
0.06 -0.3
-99.99 -0.4
-99.99 -0.5
-99.99 -0.6
-99.99 -0.5

### 24 Mar 2026 v8

# Discrete Chi-Square Method can model and forecast complex time series, like El Nino data between 1870 and 2026

@@ Abstract
Forecasting El Niño is one of the greatest challenges of science. We show how intensive, large and accurate time series allow us to see through time. Our Discrete Chi-square Method (DCM) can detect arbitrary trend and signal(-s) combinations. It can forecast complex time series. The widely-used Discrete Fourier Transform (DFT) and other frequency-domain parametric time series analysis methods have many application limitations. None of those limitations constrains the DCM. Our simulated time series analyses ascertain the revolutionary Window Dimension Effect (WD-effect): “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.” The simulations also expose the DFT’s weaknesses and the DCM’s efficiency. The DCM’s backbone is the Gauß-Markov theorem that the Least Squares (LS) is the best unbiased estimator for linear regression models. DCM can not fail because this simple method is based on the computation of a massive number of linear model LS fits. The Fisher-test gives the signal significance estimates and identifies the best DCM model from all alternative tested DCM models. The analytical solution for the non-linear DCM model is an ill-posed problem. We present a computational well-posed solution. The best DCM model must be correct if it passes our Forecast-test. Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series. In our appendix, we show that the DCM can model and forecast El Nino data between 1870 and 2026. An immediate, independent and objective validity check of our analysis may save some money.
@@ 1 Introduction
This study proceeds through following stages. We present the DCM formulation (Section 2.1). The WD-effect is discussed (Section 2.2). Complex time series are simulated using the DCM model (Sect. 2.3). The DFT and its ALs are presented (Section 2.4). We compare our DCM to the renowned DFT. The time series analysis of simulated data sets shows that the ALs of DFT do not constrain DCM (Sections 3.1-3.8). Then, we present the identification of the best DCM model for the data (Section 3.9), the DCM significance estimates (Section 3.10), the well-posed computational DCM model solution (Section 3.11) and the DCM forecasts (Section 3.12). Finally, we discuss our results (Section 4). In our Appendix, we show that the DCM can model and forecast El Niño (Section Appendix 1).
@@ 2.1 Discrete Chi-square method (DCM)
The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [32, TSPA, Equation 1] and the Continuous Period Search [40, CPS, Equation 3]. TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$. TSPA can detect pure sine and double wave signals ($K_{2}=1$ or 2). The extension of TSPA, the CPS method, tests three alternatives: $K_{2}=1$ or 2, or no signal at all. The DCM model is more sophisticated because it can have any arbitrary $K_{1}$, $K_{2}$ and $K_{3}$ combination. The DCM sum $g(t)$ =h(t)+p(t)=h(t)+p(t) of arbitrary periodic and aperiodic functions represents a universal model because the innumerable $K_{1}$, $K_{2}$ and $K_{3}$ combinations allow unlimited complexity. Even more complex signal shape $(K_{2})$ combinations could be used. For example, these shapes could be $K_{2,i=1}=3$ (3rd harmonic), $K_{2,i=2}=1$ (pure sine) and $K_{2,i=3}=2$ (double wave), where if $K_{2,i}$ denotes the shape of $i$:th signal.
@@ 2.4 Discrete Fourier Transform (DFT)
The DFT is one of the most prevalent frequency-domain parametric time series analysis methods for unevenly spaced time series. It searches for the best pure sine model for the data. The equivalent DCM model has the orders $K_{1}=1,K_{2}=1$ and $K_{3}=0$. Any time series analysis method must be remarkable if it performs better than the distinguished DFT. Therefore, we will compare how DCM and DFT perform in the analyses of simulated time series (Sections 3.1-3.7). We search for signals in these simulated time series by using the frequently applied DFT version formulated by [25] [^1] This paper had over 2600 citations in December 2025. Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [25, Equation 1]. The notations for the DFT model are
@@
The time series is “too short” [33, AL6]
@@
The time series is non-stationary because it contains “a trend” [46, AL7]
@@
Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [33, AL9].
@@ 3.9 Best model
In the above example, the Fisher-test finds the correct number of signals $(K_{1})$ and the correct trend $(K_{3})$ for the pure sine signal alternative $(K_{2}=1)$. We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large. One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu [31, Table S7].
@@ 3.10 Significance estimates
After the detection of the first strongest signal, the Fisher-test critical level $Q_{\mathrm{F}}$ values increase for the next detected weaker signals. In other words, the signal significances decrease. Typical examples can be found in Jetsu [31, Tables S5-S1]. No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected. For the Model 1 simulated $n=100$ and ${\mathrm{SN}}=100$ combination, Table 8 shows that this time series contains only one signal.
For this Model 1 time series, there is no need to discuss the DFT significance estimates [25, their Equation 22] because this method fails to detect the correct period.
@@ 3.11.3 Stability ( $\mathrm{C}_{3}$)
The artificial bootstrap data sets (Equation 31) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation). We routinely check the stability of these solutions [31, Figure S5]. The unstable models, where the model parameter changes are large, are rejected (Section 2.1: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}” models).
@@ 3.12 Forecast
There are numerous techniques for forecasting a time series [e.g., 24, 29, 34]. The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast. We divide all data into the forecasting data and the forecasted data. The time points, the observations and the errors of these samples are
@@ Results
We use the first half of all yearly data $(n=78)$ to forecast the second half $(n=77)$. The results of weighted DCM analysis are given in Table S2. The periods of the three strongest signals are $P_{1}=5.580\pm 0.085$, $P_{2}=12.82\pm 0.40$ and $P_{3}=19.30\pm 0.83$ years (Model $\mathcal{M}$=3). The forecasts of the two and three signals models, open circles, are shown in Figures S1 and S2. Both forecasts reproduce the overall yearly mean changes during the second half. For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signal models are 421, 336, 370 and 428, respectively. Hence, the two signal $\mathcal{M}$=2 model gives the best “Big Wave” forecast.
@@
For all weighted data, the detection of the first “Big Wave” signal is absolutely certain. The Fisher-test comparison between the $K_{3}=0$ pure noise model $g_{0,1,0,\chi^{2}}$ and the one signal model $g_{1,1,1,\chi^{2}}$ gives signal significance $Q_{\mathrm{F}}=3.9\times 10^{-13}$ (Table S6, $\mathcal{M}$=1). This is the probability that the first detected signal is not real. All other polynomial model alternatives, $1\leq K_{3}\leq 7$, are also rejected. The second and third “Big Wave” signal significances are $Q_{\mathrm{F}}=2.7\times 10^{-7}$ and $Q_{\mathrm{F}}=0.00038$ (Table S3: $\mathcal{M}$=2-3).
@@
Our mathematical model can detect regularities that the current physical models fail to detect, let alone forecast. Mathematics detects signals that should not be there according to Physics. The counterargument is that simple mathematics can not explain a complex physical phenomenon. We argue that the El Niño physical model for the “Big Wave” is simple. The solar energy output changes could determine the “Big Wave”. It is no surprise that the Sun warms the largest battery on the Earth, the Pacific Ocean. And if many cycles moderate the solar energy output, it is no surprise either that we have failed to discover the interference of these cycles, the “Big Wave”, in our yearly climate temperatures. The immense Pacific Ocean surface may act like a giant thermometer that measures solar energy output in real time.
@@
Appendix 2
The analysed data in Appendix 1 were the yearly values before December 2024.
Future data
We write this Appendix 2 because our research has arrived at a "Mexican standoff, where only future data can confirm our “Big Wave” forecasts. Fortunately, one minor detail saves us one year of waiting.
Year 2025
On January 29th, 2026, we computed the yearly values from the NOAA1 data. These yearly data ended in December 2024 because the four last monthly values for the year 2025 were missing. On February 25th, 2026, NOAA1 still had these same four missing monthly values for the year 2025 (“-99.99”). To get these missing monthly values, we were advised to use the Climate Prediction Center (CPC) data in NOAA2.666https://www.cpc-ncep.noaa.gov/products/
@@
All twelve monthly values for the year 2025 are given in Table S7 (column 3). These monthly values give the 2025 yearly value $y_{i}\pm 1\sigma_{i}=-0.26\pm 0.21$ $\mathrm{C}^{o}$. A red circle now highlights this observed new 2025 yearly value in Figures S3e and S4e. Our “Big Wave” forecast for the year 2025 succeeds. This 2025 forecast for the best all yearly data $\mathcal{M}$=3 model (Table S3) is more accurate than the forecast for the unstable $\mathcal{M}$=4 model (Table S3). Our forecast for this new 2025 observation can, of course, be dismissed just as one lucky guess. Who wants to take the risk that our severe “Big Wave” forecasts for the years 2030 and 2031 are also “lucky”?
All months
We did not analyse the monthly data in Appendix 1. The analysed yearly data ended in December 2024. Here, we analyse the NOAA1 monthly data ending in December 2024. This ensures that the same numbers are used to compute the yearly and the monthly data values. The computation of monthly means eliminated the detection of shorter periods from yearly data. Therefore we lower the $P_{\mathrm{min}}$ value from 3 to 2 years when analysing these monthly data. The value $P_{\mathrm{min}}=100$ years remains the same.
The DCM analysis results for all monthly values between January 1870 and December 2024 are given in Table S8. The four signal $\mathcal{M}$=4 model is the best. The signal periods are $P_{1}=3.643\pm 0.034$ years, $P_{2}=4.7475\pm 0.0082$ years $P_{3}=5.649\pm 0.012$ years and $P_{4}=12.706\pm 0.050$ years. The 12.7 year signal has the highest amplitude. The periodogram shows four clear minima (Figure S5a). The 1st signal significance is $Q_{\mathrm{F}}<\!10^{-16}$ (Table S9). The 2nd, 3rd, and 4th signal significances are also $Q_{\mathrm{F}}<\!10^{-16}$ (Table S8). These signal detections are absolutely certain. The DCM search for the fifth signal from the $\mathcal{M}$=4 model residuals gives $R=528.7$. Hence, these monthly data contain only four signals because the $\mathcal{M}$=4 model has $R=475.6$.
The best DCM model for all monthly data is shown in Figures S5b-e. The residuals show strong correlation, which is typical for climatological time series [64, 17]. Due to this correlation, the bootstrap procedure gives too small model error $\sigma_{g(t)}$ estimates (Equation 30). Therefore, only the next decade monthly forecast values, but not their error estimates, are given in Table S10.
Separate months
We also analyse all months separately because that may eliminate the residual correlation problem. The results are given in Tables S11-S22. The forecasted monthly values are compared to the NOAA2 values (Table S7, column 3), which are highlighted as red dots in Figures S6-S17. Except for November 2025, all other red dots are within the forecast error $\pm 3\sigma_{g(t)}$ red dotted lines. For separate months analysis, all next decade forecasted monthly values are given in Table S23. Note that this table also gives the forecast error estimates because residual correlation for separate months is weaker, or sometimes even totally absent.
Connection to solar cycle periods
[…]
$P_{\mathrm{Syn}}=[P_{1}^{-1}\pm P_{2}^{-1}]^{-1}$ (A1)
predicts the synodic periods of these signals given in Table S24. For example, the strongest 12.82 years period in El Niño data could arise from combination $P_{1}=11.86/2$ years and $P_{2}=11$ years given in line 15 at column 4. We think that it is not necessary to discuss all possible synodic period combinations in this short appendix.
Physics
We use the relation
$F_{\mathrm{Sun}}(t)\longrightarrow T_{\mathrm{Sea}}(t)$ (A2)
to describe how the solar output $F_{\mathrm{Sun}}]={\mathrm{W}}{\mathrm{m}}^{-2}$ triggers the Pacific Ocean sea surface temperature $T_{\mathrm{Sea}}]={\mathrm{C}}^{\mathrm{o}}$ changes. Our symbol $\longrightarrow$” contains all unknown climatological processes.
The monthly data residuals consistently show how rapidly and strongly the $T_{\mathrm{Sea}}$ values react to the $F_{\mathrm{Sun}}$ peaks and valleys (Figures S5b-e, blue dots). The $F_{\mathrm{Sun}}$ and $T_{\mathrm{Sea}}$ extrema coincide. The DCM detects the $F_{\mathrm{Sun}}$ signals. The solar maximum effect is $A_{1}+A_{2}+A_{3}+A_{4}=1.4$$\mathrm{C}^{o}$ (Table S8). However, the chaotic $\longrightarrow$” trigger effect can reach $\pm 1$$\mathrm{C}^{o}$. We can not forecast the exact $T_{\mathrm{Sea}}$ values, but we can show that the general weather pattern follows the solar energy output “Big Wave”. This does not fully solve the El Niño riddle, but it gives a good starting point for modelling the climatological $\longrightarrow$” effects. If no internal mechanism can cause the strictly multi-periodic “Big Wave” in the Pacific Ocean then the Sun is the only possible external cause. The planets may indeed cause regular solar cycles [31]. NOAA and other official agencies report El Niño probabilities like 60%, …, 90%. The DCM probability that the four strongest “Big wave” signals are real is larger than $1-10^{-16}$ (Table S8).
@@ Acknowledgements
All manuscript data files and Python codes are stored to
@@
The newest additional Appendix data files and control files are stored to
https://zenodo.org/records/19183791/
@@
Figure S3: “Big Wave” for all yearly mean data model $\mathcal{M}$=3 in Table S3. Red circle is added new observed 2025 year value. Otherwise as in Figure S1.
@@
Table S2: Periods in first half yearly mean data. Otherwise as in Table S1.
@@
Table S4: “Big Wave”yearly forecasts (Figures S3e and S4e: open circles).
@@
Table S7: Data.
(1) Month
(2) NOAA1 on 28.01.2026. (3) NOAA2 on 13.03.2026.
Month $y$ $y$
(-) $\mathrm{C}^{o}$ $\mathrm{C}^{o}$
(1) (2) (3)
Jan-25 -0.50 -0.4
Feb-25 -0.36 -0.2
Mar-25 -0.13 -0.1
Apr-25 -0.09 0.0
May-25 0.02 0.0
Jun-25 0.14 0.0
Jul-25 0.05 -0.1
Aug-25 0.06 -0.3
Sep-25 -99.99 -0.4
Oct-25 -99.99 -0.5
Nov-25 -99.99 -0.6
Dec-25 -99.99 -0.5
Jan-26 - -0.4
Feb-26 - -
Refer to caption Refer to caption
            (b)             (c)
Refer to caption Refer to caption
            (d)             (e)
Figure S5: All monthly data model $\mathcal{M}$=4 in Table S8. Red dots are monthly values from Table S7. Otherwise as in Figure S2.
Table S8: Periods in non-weighted all months data. Otherwise as in Table S1.
Period analysis Fisher-test
Data: Original non-weighted data ($n=1860,\Delta T=154.9$: mAll.dat)
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
One signal
$\mathcal{M}$=1 $12.716\pm 0.047$ - - - $\uparrow$ $\uparrow$ $\uparrow$
$\mathcal{M}$1,1,1,R 5 $0.396\pm 0.029$ - - - 32.9 33.2 35.3 mAllR14K111.dat
556.8 $1871.62\pm 0.37$ - - - $Q_{\mathrm{F}}<\!10^{-16}$ $Q_{\mathrm{F}}<\!10^{-16}$ $Q_{\mathrm{F}}<\!10^{-16}$
Two signals
$\mathcal{M}$=2 $5.651\pm 0.012$ $12.700\pm 0.042$ - - - $\uparrow$ $\uparrow$
$\mathcal{M}$2,1,1,R 8 $0.348\pm 0.029$ $0.389\pm 0.028$ - - - 31.8 34.7 mAllR14K211.dat
528.6 $1870.75\pm 0.19$ $1871.76\pm 0.32$ - - - $Q_{\mathrm{F}}<\!10^{-16}$ $Q_{\mathrm{F}}<\!10^{-16}$
Three signals
$\mathcal{M}$=3 $4.7459\pm 0.0082$ $5.640\pm 0.012$ $12.694\pm 0.043$ - - - $\uparrow$
$\mathcal{M}$3,1,1,R 11 $0.335\pm 0.023$ $0.340\pm 0.034$ $0.388\pm 0.030$ - - - 35.7 mAllR14K311.dat
502.6 $1870.60\pm 0.13$ $1870.93\pm 0.15$ $1871.83\pm 0.29$ - - - $Q_{\mathrm{F}}<\!10^{-16}$
Four signals
$\mathcal{M}$=4 $3.6436\pm 0.0046$ $4.7477\pm 0.0080$ $5.647\pm 0.011$ $12.704\pm 0.063$ - - -
$\mathcal{M}$4,1,1,R 14 $0.345\pm 0.030$ $0.332\pm 0.031$ $0.342\pm 0.025$ $0.392\pm 0.038$ - - - mAllR14K411.dat
475.0 $1872.68\pm 0.11$ $1870.55\pm 0.14$ $1870.82\pm 0.19$ $1871.74\pm 0.39.$ - - -
@@
$g_{1,1,1}$
Polynomial $\eta=5,R=557.1$
$\mathcal{M}$=1 $\uparrow$
$g(t)=p(t,K_{3}=0)$ $F=63.9$
$\eta=1$, $R=634$ $Q_{\mathrm{F}}<\!10^{-16}$
$\mathcal{M}$=2 $\uparrow$
$g(t)=p(t,K_{3}=1)$ $F=39.8$
$\eta=2$, $R=593$ $Q_{\mathrm{F}}<\!10^{-16}$
$\mathcal{M}$=3 $\uparrow$
$g(t)=p(t,K_{3}=2)$ $F=59.7$
$\eta=3$, $R=593$ $Q_{\mathrm{F}}<\!10^{-16}$
$\mathcal{M}$=4 $\uparrow$
$g(t)=p(t,K_{3}=3)$ $F=92.8$
$\eta=4$, $R=585$ $Q_{\mathrm{F}}<\!10^{-16}$
$\mathcal{M}$=5 $\uparrow$
$g(t)=p(t,K_{3}=4)$ No test11footnotemark: 1
$\eta=5$, $R=584$ $Q_{\mathrm{F}}=1$
$\mathcal{M}$=6 $\uparrow$
$g(t)=p(t,K_{3}=5)$ F=34F=-34
$\eta=6$, $R=584$ $Q_{\mathrm{F}}=1$
$\mathcal{M}$=7 $\uparrow$
$g(t)=p(t,K_{3}=6)$ F=33.5F=-33.5
$\eta=7$, $R=578$ $Q_{\mathrm{F}}=1$
$\mathcal{M}$=8 $\uparrow$
$g(t)=p(t,K_{3}=7)$ F=22.3F=-22.3
$\eta=8$, $R=578$ $Q_{\mathrm{F}}=1$
[…]
Table S10: All months. “Big Wave” forecast (Figure S5). Symbols $\uparrow$ and $\downarrow$ tell that preceding value is smaller or bigger.
Year Jan Feb Mar Apr May Jun Jul Aug Sep Oct Nov Dec
2025 0.06 $\uparrow$ 0.05 $\downarrow$ 0.04 $\downarrow$ 0.03 $\downarrow$ 0.02 $\downarrow$ 0.02 $\downarrow$ 0.01 $\downarrow$ 0.00 $\downarrow$ -0.00 $\downarrow$ -0.01 $\downarrow$ -0.01 $\downarrow$ -0.01 $\uparrow$
2026 -0.00 $\uparrow$ 0.00 $\uparrow$ 0.01 $\uparrow$ 0.01 $\uparrow$ 0.02 $\uparrow$ 0.03 $\uparrow$ 0.04 $\uparrow$ 0.05 $\uparrow$ 0.06 $\uparrow$ 0.08 $\uparrow$ 0.09 $\uparrow$ 0.10 $\uparrow$
2027 0.11 $\uparrow$ 0.12 $\uparrow$ 0.12 $\uparrow$ 0.13 $\uparrow$ 0.13 $\uparrow$ 0.13 $\uparrow$ 0.13 $\uparrow$ 0.13 $\downarrow$ 0.13 $\downarrow$ 0.12 $\downarrow$ 0.12 $\downarrow$ 0.11 $\downarrow$
2028 0.10 $\downarrow$ 0.09 $\downarrow$ 0.09 $\downarrow$ 0.08 $\downarrow$ 0.07 $\downarrow$ 0.06 $\downarrow$ 0.05 $\downarrow$ 0.05 $\downarrow$ 0.04 $\downarrow$ 0.04 $\downarrow$ 0.04 $\uparrow$ 0.04 $\uparrow$
2029 0.05 $\uparrow$ 0.06 $\uparrow$ 0.07 $\uparrow$ 0.08 $\uparrow$ 0.10 $\uparrow$ 0.12 $\uparrow$ 0.14 $\uparrow$ 0.17 $\uparrow$ 0.19 $\uparrow$ 0.22 $\uparrow$ 0.25 $\uparrow$ 0.28 $\uparrow$
2030 0.31 $\uparrow$ 0.34 $\uparrow$ 0.37 $\uparrow$ 0.40 $\uparrow$ 0.42 $\uparrow$ 0.45 $\uparrow$ 0.47 $\uparrow$ 0.49 $\uparrow$ 0.50 $\uparrow$ 0.52 $\uparrow$ 0.52 $\uparrow$ 0.53 $\uparrow$
2031 0.53 $\uparrow$ 0.52 $\downarrow$ 0.52 $\downarrow$ 0.50 $\downarrow$ 0.49 $\downarrow$ 0.47 $\downarrow$ 0.45 $\downarrow$ 0.42 $\downarrow$ 0.39 $\downarrow$ 0.36 $\downarrow$ 0.33 $\downarrow$ 0.30 $\downarrow$
2032 0.27 $\downarrow$ 0.23 $\downarrow$ 0.20 $\downarrow$ 0.17 $\downarrow$ 0.14 $\downarrow$ 0.12 $\downarrow$ 0.09 $\downarrow$ 0.07 $\downarrow$ 0.05 $\downarrow$ 0.03 $\downarrow$ 0.02 $\downarrow$ 0.01 $\downarrow$
2033 0.01 $\downarrow$ 0.01 $\downarrow$ 0.01 $\uparrow$ 0.02 $\uparrow$ 0.02 $\uparrow$ 0.03 $\uparrow$ 0.05 $\uparrow$ 0.06 $\uparrow$ 0.08 $\uparrow$ 0.09 $\uparrow$ 0.11 $\uparrow$ 0.13 $\uparrow$
2034 0.14 $\uparrow$ 0.15 $\uparrow$ 0.17 $\uparrow$ 0.17 $\uparrow$ 0.18 $\uparrow$ 0.19 $\uparrow$ 0.19 $\uparrow$ 0.18 $\downarrow$ 0.18 $\downarrow$ 0.16 $\downarrow$ 0.15 $\downarrow$ 0.13 $\downarrow$
2035 0.11 $\downarrow$ 0.09 $\downarrow$ 0.06 $\downarrow$ 0.03 $\downarrow$ 0.00 $\downarrow$ -0.03 $\downarrow$ -0.06 $\downarrow$ -0.10 $\downarrow$ -0.13 $\downarrow$ -0.16 $\downarrow$ -0.19 $\downarrow$ -0.22 $\downarrow$